OpenStudy (anonymous):

HELPPPPPPPP write 5x-2y=3 in point slope intercept form

OpenStudy (anonymous):

is it Y-3/2 = 5/2 (x-2)?

OpenStudy (anonymous):

medal

OpenStudy (anonymous):

yes i have that and then i did y-3/2+5/2x and then im having trouble at the end

TheSmartOne (thesmartone):

5/2 would be the correct slope. And when x=0, y will = to -3/2

TheSmartOne (thesmartone):

So y+3/2=5/2(x-0)

TheSmartOne (thesmartone):

@jleonardo12

OpenStudy (anonymous):

yeah because the point is (0,3/2) but the choices are Y-3/2 = 5/2 (x-2) or y-3/2+5/2(x-3/2) which makes no sense

TheSmartOne (thesmartone):

Are you sure you wrote the question correctly..

TheSmartOne (thesmartone):

@jleonardo12

OpenStudy (anonymous):

yes it might be a typo but I'm taking an online test so i don't have a professor to ask if it is